Can someone help me on this problem?
I need to solve the following word problem. I also need to show the equation you use for the solution.
A bicyclist rode into the country for 5 h. In returning, her speed was 5 mi/h faster and the trip took 4 h. What was her speed each way?
Thanks,
Ashley

Answer by stanbon(75887)

A bicyclist rode into the country for 5 h. In returning, her speed was 5 mi/h faster and the trip took 4 h. What was her speed each way?
--------
Riding-Out DATA:
time=5 hr. ; rate = x mph ; distance=5x miles
---------
Riding-Back DATA:
time=4 hr. ; rate= x+5 mph ; distance=rt = 4(x+5)=4x+20 miles
-----------
EQUATION:
distace out = distance back
5x = 4x+20
x=20 mph (speed on the way out)
x+5=25 mph (speed on the way back)
Cheers,
Stan H.